OpenStudy (anonymous):

Solve the equation for x and round your answer to the nearest hundredth. 5 = ln(x + 1)

OpenStudy (unklerhaukus):

\[5 = \ln(x + 1)\] first step is to raise both sides of the equation to the power e \[e^5 = e^{\ln(x + 1)}=x+1\]

OpenStudy (unklerhaukus):

now, to solve for x , just take one away from both sides

OpenStudy (unklerhaukus):

do you understand what i have said?

OpenStudy (anonymous):

yes thank you